a block of wood and a 0.90 kg block of steel are placed in thermal contact while thermally isolated from their surroundings of the the wood was at an initial temperature of 40 degrees the steel was at an initial temperature of 60 degrees and the final equilibrium temperature of the wood and steel was 45 degrees then what was the mass of the block of wood


1
Expert's answer
2021-05-27T12:30:04-0400

Gives

"M_1=0.90kg"

"T_1=40\u00b0C"

"T_2=60\u00b0C"

"T=45\u00b0C"

"M_2=?"

We know that

"T=\frac{M_1T_1+M_2T_2}{M_1+M_2}"

Put value

"0.90+M_2=\frac{0.90\times40+M_2\times60}{45}"


"(0.90+M_2)\times45=0.90\times45-0.90\times40"

"M_2(60-45)=0.90(45-40)"

"M_2=\frac{0.90}{3}=0.30kg"

"M_2=0.30 kg"